jhonyy9 (jhonyy9):

given : 1 + 4 = 5 2 + 5 = 12 3 + 6 = 21 8 + 11 = ?

OpenStudy (lowkey.s):

Okay.

jhonyy9 (jhonyy9):

so really i m not sure but said somebody that are ,exist two solutions my opinion that just 96 is the true,really possible number there on place of question mark what is your opinion please ???? ty. @agent0smith @TheSmartOne @ganeshie8

OpenStudy (lowkey.s):

1 + 4 = 5 5 + 2 + 5 = 12 12 + 3 + 6 = 21 21 + 8 + 11 = 40

OpenStudy (lowkey.s):

1 + (1 x 4) = 5 2 + (2 x 5) = 12 3 + (3 x 6) = 21 8 + (8 x 11) = 96 or

OpenStudy (lowkey.s):

either 40 or 96

jhonyy9 (jhonyy9):

this is the second what have wrote me on facebook.com but this 40 is true too than i will rewrite these equations in this line 1 + 4 = 5 3 + 6 = 18 8 + 11 = ? 5 + 2 = 12 so in this case what will be the right answer ?

jhonyy9 (jhonyy9):

sorry there is 2 + 5 = 12 the last one line

jhonyy9 (jhonyy9):

so in this order wrote these equations can being 40 true again too ? i think not

jhonyy9 (jhonyy9):

in this case just 96 is the right answer

OpenStudy (lowkey.s):

No Its not 40

OpenStudy (loser66):

I think only 96 is the correct answer

jhonyy9 (jhonyy9):

Loser agent

OpenStudy (sshayer):



OpenStudy (lowkey.s):

i argee with loser

OpenStudy (loser66):

I use "second difference" to find the logic of the sequence and get 96

OpenStudy (loser66):

And it is not a guess!! (lol)

jhonyy9 (jhonyy9):

ty.